The area of a triangle with height 5 inches is equal to the area of a certain rectangle of length 10 inches and width 7 inches. What is the base of the triangle?
Answer by mananth(16949)

length of rectangle = 10
width of rectangle =7
Area = 70 sq.in
..
Area of triangle = 1/2 * base * h
h = 5 in
Area =70
base = area *2 / h
base = 70*2/5
=14*2
= 28 inches height